# On a class of thin obstacle-type problems for the bi-Laplacian operator

### v2

@@ 6. Almgren’s blow-up solutions
We recall that, in view of the local character of the analysis we employed, the conclusions of the previous section continue to hold for (sufficiently small) balls centered at $(x_{0},0)$, for any $x_{0}\in B^{\prime}_{1}$. Without loss of generality, throughout this section we continue to assume that $x_{0}=0$.
@@
$u_{r}(X)\coloneqq\frac{u(rX)}{h(r)^{1/2}}\qquad\text{ and }\qquad v_{r}(X)\coloneqq\frac{v(rX)}{h(r)^{1/2}}.$ (6.1)
@@
$u_{r_{j}}\to\tilde{u},\quad v_{r_{j}}$ $\to\tilde{v}\qquad\text{ in }H^{1}(B_{R}^{+};y^{b}),$ (6.2)
$u_{r_{j}}\to\tilde{u},\quad v_{r_{j}}$ $\to\tilde{v}\qquad\text{ in }C^{1,\alpha}(B_{R}^{+}),$ (6.3)
as $j\to\infty$, for some $\alpha\in(0,1)$. Moreover, $\mu\in\mathbb{N}\cup\{0\}$, and $\tilde{u}$ and $\tilde{v}$ are homogeneous polynomial of degree $\mu$ that cannot simultaneously be identically equal to zero and satisfy
@@
$\int_{B_{R_{0}}^{+}}y^{b}(|\nabla u_{r}|^{2}+|\nabla v_{r}|^{2})\,dX=\frac{1}{R_{0}}N_{0}(rR_{0},u,v)H(R_{0},u_{r},v_{r})\leq(\mu+1)R_{0}^{n+b+2\mu}.$ (6.7)
@@
This shows that the families $\{u_{r}\}_{r}$ and $\{v_{r}\}_{r}$ are bounded in $H^{1}(B_{R_{0}}^{+};y^{b})$, provided that $r$ is sufficiently small. In particular, for every sequence $\{r_{j}\}_{j}$, $j\geq 2$, such that $r_{j}\to 0^{+}$, there exist a subsequence (which we do not relabel) and functions $\tilde{u},\tilde{v}\in H^{1}(B_{R_{0}}^{+};y^{b})$ such that $u_{r_{j}}\to\tilde{u}$ and $v_{r_{j}}\to\tilde{v}$ strongly in $L^{2}(B_{R_{0}}^{+};y^{b})$ (see, for example, [32]), strongly in $L^{2}(S_{1}^{+};y^{b})$ (see ˜2.2), and weakly in $H^{1}(B_{R_{0}}^{+};y^{b})$. Throughout the remainder of the proof, we continue to work with this subsequence; however, we choose to omit the subscript $j$ to keep the notation as simple as possible. In particular, the convergence of the traces obtained above implies that
$\int_{S_{1}^{+}}y^{b}(\tilde{u}^{2}+\tilde{v}^{2})\,d\mathcal{H}^{n}=\lim_{r\to 0^{+}}\int_{S_{1}^{+}}y^{b}(u_{r}^{2}+v_{r}^{2})\,d\mathcal{H}^{n}=1,$ (6.8)
@@
$\|v_{r}\|_{L^{q}(B_{R_{1}}^{+})}+\|\tilde{f}(\cdot,u_{r}(\cdot,0))\|_{L^{q}(B_{R_{1}}^{\prime})}\leq C.$ (6.10)
@@
$\left|\int_{B_{R}^{+}}y^{b}\nabla v_{r}\cdot\nabla\phi\,dX\right|\leq\int_{B_{R}^{\prime}}|\tilde{f}(x,u_{r}(x,0))\phi(x,0)|\,dx\leq Cr^{1-b}\int_{B_{R}^{\prime}}|u_{r}(x,0)|^{p-1}|\phi(x,0)|\,dx.$ (6.11)
@@
must hold for every $\phi\in H^{1}_{0,S_{R}^{+}}$. In particular, $\tilde{u}$ and $\tilde{v}$ are smooth by ˜2.10 $(i)$. Next, observe that arguing as in (6.5), if $\rho\leq R$ we have that
@@
where in the last equality we have used (6.2). In particular, this implies that $\frac{d}{d\rho}N_{0}(\rho,\tilde{u},\tilde{v})=0$ for $\mathcal{L}^{1}$-a.e. $\rho<R$. In turn, by adapting the computations in Step 1 of ˜5.6 to $\tilde{u}$ and $\tilde{v}$, we obtain that
$\left(\int_{S_{\rho}^{+}}y^{b}(\tilde{u}^{2}+\tilde{v}^{2})\,d\mathcal{H}^{n}\right)\left(\int_{S_{\rho}^{+}}y^{b}(\tilde{u}_{\nu}^{2}+\tilde{v}_{\nu}^{2})\,d\mathcal{H}^{n}\right)-\left(\int_{S_{\rho}^{+}}y^{b}(\tilde{u}\tilde{u}_{\nu}+\tilde{v}\tilde{v}_{\nu})\right)^{2}\,d\mathcal{H}^{n}=0.$ (6.12)
Since (6.12) corresponds to the equality case in the Cauchy-Schwartz inequality, the desired result follows as in the proof of Lemma 5.1 in [23]. This concludes the proof. ∎
@@
In view of ˜6.1, in the sequel it will be convenient to denote with $\mathcal{P}_{\mu}$ the space of consisting of all $\mathcal{L}_{b}$-harmonic polynomials that are homogeneous of degree $\mu$ (with $\mu\in\mathbb{N}_{0}$) and symmetric with respect to the hyperplane $\{y=0\}$. As one can readily check, every $w\in\mathcal{P}_{\mu}$ satisfies
@@ 7. Monotonicity formulas and homogeneous blow-up solutions
$M_{\mu}(r,u,v,p,q)\coloneqq\frac{1}{r^{n+b+2\mu}}\int_{S_{r}^{+}}y^{b}\left((u-p)^{2}+(v-q)^{2}\right)\,d\mathcal{H}^{n},$ (7.1)
@@
$\frac{r}{2}\frac{d}{dr}M_{\mu}(r,u,v,p,q)=\frac{1}{r^{n+b+2\mu}}(rD(r,w_{1},w_{2})-\mu H(r,w_{1},w_{2})).$ (7.5)
@@
$\frac{r}{2}\frac{d}{dr}M_{\mu}(r,u,v,p,q)=W_{\mu}(r,u,v)+\frac{rP(r,w_{1},w_{2})}{r^{n+b+2\mu}}.$ (7.6)
@@
$N_{0}(r)-\mu\geq-\frac{Cr(1+\mu)^{2}}{1+Cr(1+\mu)}\geq-Cr.$
Consequently, applying ˜5.8 $(i)$, we obtain
@@
$M_{\mu}(r_{j},u,v,\tilde{u},\tilde{v})\leq\frac{h(r_{j})}{r_{j}^{2\mu}}-\frac{2h(r_{j})^{1/2}}{r_{j}^{\mu}}+1.$
Moreover, letting $j\to\infty$, it follows from (7.9) that
$M_{\mu}(0^{+},u,v,\tilde{u},\tilde{v})=\lim_{j\to\infty}M_{\mu}(r_{j},u,v,\tilde{u},\tilde{v})=\int_{S_{1}^{+}}y^{b}(\tilde{u}^{2}+\tilde{v}^{2})\,d\mathcal{H}^{n}.$ (7.10)
@@
$Cr\leq M_{\mu}(r,u,v,\tilde{u},\tilde{v})-M_{\mu}(0^{+},u,v,\tilde{u},\tilde{v})=\frac{h(r)}{r^{2\mu}}-\frac{2h(r)^{1/2}}{r^{\mu}}\int_{S_{1}^{+}}y^{b}(u_{r}\tilde{u}+v_{r}\tilde{v})\,d\mathcal{H}^{n},$ (7.11)
@@
$Cr_{j}^{1-\delta}\leq\frac{h(r_{j})^{1/2}}{r_{j}^{\mu}}-2\int_{S_{1}^{+}}y^{b}(u_{r_{j}}\tilde{u}+v_{r_{j}}\tilde{v})\,d\mathcal{H}^{n}.$
@@
$u_{\mu,r}(X)\coloneqq\frac{u(rX)}{r^{\mu}},\qquad v_{\mu,r}(X)\coloneqq\frac{v(rX)}{r^{\mu}}.$ (7.12)
@@
$u_{\mu,r}\to\tilde{u}_{\mu},\quad v_{\mu,r}$ $\to\tilde{v}_{\mu}\qquad\text{ in }H^{1}(B_{R}^{+};y^{b}),$ (7.13)
$u_{\mu,r}\to\tilde{u}_{\mu},\quad v_{\mu,r}$ $\to\tilde{v}_{\mu}\qquad\text{ in }C^{1,\alpha}(B_{R}^{+}),$ (7.14)
for some $\alpha\in(0,1)$ as $r\to 0^{+}$.
@@
we readily obtain that (7.13) and (7.14) hold for $r=r_{j}$.
@@
$=\lim_{j\to\infty}M_{\mu}(1,u_{\mu,r_{j}},v_{\mu,r_{j}},\tilde{u}_{\mu},\tilde{v}_{\mu})=0,$ (7.16)
@@
$M_{\mu}(0^{+},u,v,\tilde{U}_{\mu},\tilde{V}_{\mu})=\lim_{j\to\infty}M_{\mu}(1,u_{\mu,R_{j}},v_{\mu,R_{j}},\tilde{U}_{\mu},\tilde{V}_{\mu})=0.$ (7.17)
@@ 8. The free boundary: regularity and structural properties
The remainder of the section focuses on the study of the singular part. Given $X_{0}\in S(u)$, we use $N_{0}^{X_{0}}(r)$ to denote the Almgren frequency functional centered at $X_{0}$ (see (5.1)); to be precise,
$N_{0}^{X_{0}}(r)\coloneqq N_{0}(r,u(X_{0}+\cdot),v(X_{0}+\cdot)).$ (8.1)
@@
With these definitions at hand, we proceed to prove that the polynomials $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ depend continuously on $X_{0}\in S_{\mu}(u)$.
@@
$M_{\mu}^{\overline{X}}(r,u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})\leq M_{\mu}^{\overline{X}}(r_{\varepsilon},u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})+C(r_{\varepsilon}-r)\leq C\varepsilon$ (8.4)
holds for all $r\in(0,r_{\varepsilon})$. In particular, this implies that
$\int_{S_{1}^{+}}y^{b}\left((\tilde{u}_{\mu}^{\overline{X}}-\tilde{u}_{\mu}^{X_{0}})^{2}+(\tilde{v}_{\mu}^{\overline{X}}-\tilde{v}_{\mu}^{X_{0}})^{2}\right)\,d\mathcal{H}^{n}=\lim_{r\to 0^{+}}M_{\mu}^{\overline{X}}(r,u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})\leq C\varepsilon.$ (8.5)
@@
$\int_{B_{1}^{+}}y^{b}\left((u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}})^{2}+(v_{\mu,r}^{\overline{X}}-\tilde{v}_{\mu}^{\overline{X}})^{2}\right)\,dX\leq C\varepsilon.$ (8.6)
@@
$\leq\int_{0}^{1}2\rho^{n+b+2\mu}\left(M_{\mu}^{\overline{X}}(\rho r,u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})+\int_{S_{1}^{+}}y^{b}\left((\tilde{u}_{\mu}^{\overline{X}}-\tilde{u}_{\mu}^{X_{0}})^{2}+(\tilde{v}_{\mu}^{\overline{X}}-\tilde{v}_{\mu}^{X_{0}})^{2}\right)\,d\mathcal{H}^{n}\right)\,d\rho.$ (8.7)
@@
$\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{C^{0,\alpha}(B_{3/4}^{+})}\leq C\left(\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{L^{2}(B_{1}^{+};y^{b})}+\|r^{2}v_{\mu,r}^{\overline{X}}\|_{L^{q}(B_{1}^{+};y^{b})}\right).$ (8.9)
@@
Finally, if $n\geq 4$, since by assumption we have that $\mu<4/(n-3)$, then we can find $q\in((2+2\mu)/2,(n+1)/2)$. This shows that there exists $C=C(\varepsilon)$, independent of $\overline{X}$ and $r$ such that $C(\varepsilon)\to 0^{+}$ as $\varepsilon\to 0^{+}$ and with the property that
$\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{L^{\infty}(B_{3/4}^{+})}\leq C(\varepsilon)$ (8.11)
@@
$\|v_{\mu,r}^{\overline{X}}-\tilde{v}_{\mu}^{\overline{X}}\|_{L^{\infty}(B_{1/2}^{+})}\leq C\left(\|v_{\mu,r}^{\overline{X}}-\tilde{v}_{\mu}^{\overline{X}}\|_{L^{2}(B_{3/4}^{+};y^{b})}+\|\bar{f}(\cdot,u_{\mu,r}^{\overline{X}})\|_{L^{\infty}(B_{3/4}^{\prime})}\right).$ (8.12)
@@
$\|\bar{f}(\cdot,u_{\mu,r}^{\overline{X}})\|_{L^{\infty}(B_{3/4}^{\prime})}\leq Cr^{1-b+\mu(p-2)}\Big(\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{L^{\infty}(B_{3/4}^{\prime})}\\ +\|\tilde{u}_{\mu}^{\overline{X}}-\tilde{u}_{\mu}^{X_{0}}\|_{L^{\infty}(B_{3/4}^{\prime})}+\|\tilde{u}_{\mu}^{X_{0}}\|_{L^{\infty}(B_{3/4}^{\prime})}\Big)^{p-1}.$ (8.14)
@@
$d_{\mu}^{X_{0}}\coloneqq\min\left\{d_{u,\mu}^{X_{0}},d_{v,\mu}^{X_{0}}\right\},$ (8.15)

### v3

@@ 6. Almgren’s blow-up solutions
We recall that, in view of the local character of the analysis we employed, the conclusions of the previous section continue to hold for (sufficiently small) balls centered at $(x_{0},0)$, for any $x_{0}\in B^{\prime}_{1}$. Throughout this section, we continue to assume $x_{0}=0$ and $b\leq 0$.
@@
(6.1) $u_{r}(X)\coloneqq\frac{u(rX)}{h(r)^{1/2}}\qquad\text{ and }\qquad v_{r}(X)\coloneqq\frac{v(rX)}{h(r)^{1/2}}.$
@@
(6.2) $u_{r_{j}}\to\tilde{u},\quad v_{r_{j}}$ $\to\tilde{v}\qquad\text{ in }H^{1}(B_{R}^{+};y^{b}),$
(6.3) $u_{r_{j}}\to\tilde{u},\quad v_{r_{j}}$ $\to\tilde{v}\qquad\text{ in }C^{1,\alpha}(B_{R}^{+}),$
as $j\to\infty$, for some $\alpha\in(0,1)$. Moreover, $\mu\in\mathbb{N}\cup\{0\}$, and $\tilde{u}$ and $\tilde{v}$ are homogeneous functions of degree $\mu$ that cannot simultaneously be identically equal to zero and satisfy
@@
(6.7) $\int_{B_{R_{0}}^{+}}y^{b}(|\nabla u_{r}|^{2}+|\nabla v_{r}|^{2})\,dX=\frac{1}{R_{0}}N_{0}(rR_{0},u,v)H(R_{0},u_{r},v_{r})\leq(\mu+1)R_{0}^{n+b+2\mu}.$
@@
This shows that the families $\{u_{r}\}_{r}$ and $\{v_{r}\}_{r}$ are bounded in $H^{1}(B_{R_{0}}^{+};y^{b})$, provided that $r$ is sufficiently small. In particular, for every sequence $\{r_{j}\}_{j}$, $j\geq 2$, such that $r_{j}\to 0^{+}$, there exist a subsequence (which we do not relabel) and functions $\tilde{u},\tilde{v}\in H^{1}(B_{R_{0}}^{+};y^{b})$ such that $u_{r_{j}}\to\tilde{u}$ and $v_{r_{j}}\to\tilde{v}$ strongly in $L^{2}(B_{R_{0}}^{+};y^{b})$ (see, for example, [33]), strongly in $L^{2}(S_{1}^{+};y^{b})$ (see ˜2.2), and weakly in $H^{1}(B_{R_{0}}^{+};y^{b})$. Throughout the remainder of the proof, we continue to work with this subsequence; however, we choose to omit the subscript $j$ to keep the notation as simple as possible. In particular, keeping in mind the definition of $h(r)$ (5.33), the convergence of the traces obtained above implies that
(6.8) $\int_{S_{1}^{+}}y^{b}(\tilde{u}^{2}+\tilde{v}^{2})\,d\mathcal{H}^{n}=\lim_{r\to 0^{+}}\int_{S_{1}^{+}}y^{b}(u_{r}^{2}+v_{r}^{2})\,d\mathcal{H}^{n}=1,$
@@
(6.10) $\|v_{r}\|_{L^{q}(B_{R_{1}}^{+})}+\|\tilde{f}(\cdot,u_{r}(\cdot,0))\|_{L^{q}(B_{R_{1}}^{\prime})}\leq C.$
@@
(6.11) $\left|\int_{B_{R}^{+}}y^{b}\nabla v_{r}\cdot\nabla\phi\,dX\right|\leq\int_{B_{R}^{\prime}}|\tilde{f}(x,u_{r}(x,0))\phi(x,0)|\,dx\leq Cr^{1-b}\int_{B_{R}^{\prime}}|u_{r}(x,0)|^{p-1}|\phi(x,0)|\,dx.$
@@
must hold for every $\phi\in H^{1}_{0,S_{R}^{+}}$. In particular, $\tilde{u}$ and $\tilde{v}$ are smooth by ˜2.10 $(i)$ and computations analogous to those in the proof of ˜5.5 yield
(6.12) $\frac{d}{d\rho}D_{0}(\rho,\tilde{u},\tilde{v})=\frac{n+b-1}{\rho}D_{0}(\rho,\tilde{u},\tilde{v})+2\int_{S_{\rho}^{+}}y^{b}(\tilde{u}_{\nu}^{2}+\tilde{v}_{\nu}^{2})\,d\mathcal{H}^{n}.$
Next, observe that arguing as in (6.5), if $\rho\leq R$ we have that
@@
where in the last equality we have used (6.2). Following the computations in Step 1 in the proof of ˜5.6), by (6.12) and (6.13) we obtain that
[…]
$=2\rho\frac{\int_{S_{\rho}^{+}}y^{b}(\tilde{u}_{\nu}^{2}+\tilde{v}_{\nu}^{2})\,d\mathcal{H}^{n}}{\int_{S_{\rho}^{+}}y^{b}(\tilde{u}^{2}+\tilde{v}^{2})\,d\mathcal{H}^{n}}-2\rho\left(\frac{\int_{S_{\rho}^{+}}y^{b}(\tilde{u}\tilde{u}_{\nu}+\tilde{v}\tilde{v}_{\nu})\,d\mathcal{H}^{n}}{\int_{S_{\rho}^{+}}y^{b}(\tilde{u}^{2}+\tilde{v}^{2})\,d\mathcal{H}^{n}}\right)^{2}.$
In particular, we have that
(6.14) $\left(\int_{S_{\rho}^{+}}y^{b}(\tilde{u}^{2}+\tilde{v}^{2})\,d\mathcal{H}^{n}\right)\left(\int_{S_{\rho}^{+}}y^{b}(\tilde{u}_{\nu}^{2}+\tilde{v}_{\nu}^{2})\,d\mathcal{H}^{n}\right)-\left(\int_{S_{\rho}^{+}}y^{b}(\tilde{u}\tilde{u}_{\nu}+\tilde{v}\tilde{v}_{\nu})\ d\mathcal{H}^{n}\right)^{2}\,=0.$
Since (6.14) corresponds to the equality case in the Cauchy-Schwarz inequality, the desired result follows as in the proof of Lemma 5.1 in [24]. This concludes the proof. ∎
@@
It is worth noting that arguing exactly as in the proof of Lemma 5.1 in [24] yields a finer description of the blow-up profiles found in ˜6.1. In particular, we have that
(6.15) $\tilde{u}(X)=|X|^{\mu}\Psi_{1}(X/|X|)\qquad\text{ and }\qquad\tilde{v}(X)=|X|^{\mu}\Psi_{2}(X/|X|),$
where $\Psi_{1}$ and $\Psi_{2}$ are linear combinations of eigenfunctions of
$\left\{\begin{array}[]{rll}-\operatorname{div}(y^{b}\nabla_{S_{1}^{+}}\Psi)=&\lambda y^{b}\Psi&\text{ in }S_{1}^{+},\\ \partial_{y}^{b}\Psi=&0&\text{ on }\partial S_{1}^{+}\end{array}\right$.
associated with the same eigenvalue $\lambda=\lambda(n,b,\mu)$. For this reason, in the sequel it will be convenient to denote by $\mathcal{P}_{\mu}$ the space consisting of all $\mathcal{L}_{b}$-harmonic functions that are symmetric with respect to the hyperplane $\{y=0\}$ and admit a representation of the form (6.15). As one can readily check, every $w\in\mathcal{P}_{\mu}$ satisfies
@@ 7. Monotonicity formulas and homogeneous blow-up solutions
(7.1) $M_{\mu}(r,u,v,p,q)\coloneqq\frac{1}{r^{n+b+2\mu}}\int_{S_{r}^{+}}y^{b}\left((u-p)^{2}+(v-q)^{2}\right)\,d\mathcal{H}^{n},$
@@
(7.5) $\frac{r}{2}\frac{d}{dr}M_{\mu}(r,u,v,p,q)=\frac{1}{r^{n+b+2\mu}}(rD(r,w_{1},w_{2})-\mu H(r,w_{1},w_{2})).$
@@
(7.6) $\frac{r}{2}\frac{d}{dr}M_{\mu}(r,u,v,p,q)=W_{\mu}(r,u,v)+\frac{rP(r,w_{1},w_{2})}{r^{n+b+2\mu}}.$
@@
$N_{0}(r)-\mu\geq-\frac{Cr(1+\mu)^{2}}{1+Cr(1+\mu)}\geq-Cr,$
possibly increasing the value of $C$ as needed. Consequently, applying ˜5.8 $(i)$, we obtain
@@
(7.10) $\frac{h(r_{j})}{r_{j}^{2\mu}}-\frac{2h(r_{j})^{1/2}}{r_{j}^{\mu}}+1\leq M_{\mu}(r_{j},u,v,\tilde{u},\tilde{v})\leq\frac{h(r_{j})}{r_{j}^{2\mu}}+\frac{2h(r_{j})^{1/2}}{r_{j}^{\mu}}+1.$
Letting $j\to\infty$ in (7.10), it follows from (7.9) that
(7.11) $M_{\mu}(0^{+},u,v,\tilde{u},\tilde{v})=\lim_{j\to\infty}M_{\mu}(r_{j},u,v,\tilde{u},\tilde{v})=1.$
@@
(7.12) $Cr\leq M_{\mu}(r,u,v,\tilde{u},\tilde{v})-M_{\mu}(0^{+},u,v,\tilde{u},\tilde{v})=\frac{h(r)}{r^{2\mu}}-\frac{2h(r)^{1/2}}{r^{\mu}}\int_{S_{1}^{+}}y^{b}(u_{r}\tilde{u}+v_{r}\tilde{v})\,d\mathcal{H}^{n},$
@@
$Cr_{j}^{1-\delta/2}\leq\frac{h(r_{j})^{1/2}}{r_{j}^{\mu}}-2\int_{S_{1}^{+}}y^{b}(u_{r_{j}}\tilde{u}+v_{r_{j}}\tilde{v})\,d\mathcal{H}^{n}.$
@@
(7.13) $u_{\mu,r}(X)\coloneqq\frac{u(rX)}{r^{\mu}},\qquad v_{\mu,r}(X)\coloneqq\frac{v(rX)}{r^{\mu}}.$
@@
(7.14) $u_{\mu,r}\to\tilde{u}_{\mu},\quad v_{\mu,r}$ $\to\tilde{v}_{\mu}\qquad\text{ in }H^{1}(B_{R}^{+};y^{b}),$
(7.15) $u_{\mu,r}\to\tilde{u}_{\mu},\quad v_{\mu,r}$ $\to\tilde{v}_{\mu}\qquad\text{ in }C^{1,\alpha}(B_{R}^{+}),$
for some $\alpha\in(0,1)$ as $r\to 0^{+}$. Additionally, $\tilde{u}_{\mu}$ and $\tilde{v}_{\mu}$ cannot simultaneously be identically equal to zero.
@@
we readily obtain that (7.14) and (7.15) hold for $r=r_{j}$. Observe that $\tilde{u}_{\mu}$ and $\tilde{v}_{\mu}$ cannot both be identically zero, since they are multiples of the homogeneous functions $\tilde{u}$ and $\tilde{v}$ obtained in ˜6.1, which themselves are nontrivial.
@@
(7.17) $=\lim_{j\to\infty}M_{\mu}(1,u_{\mu,r_{j}},v_{\mu,r_{j}},\tilde{u}_{\mu},\tilde{v}_{\mu})=0,$
@@
(7.18) $M_{\mu}(0^{+},u,v,\tilde{U}_{\mu},\tilde{V}_{\mu})=\lim_{j\to\infty}M_{\mu}(1,u_{\mu,R_{j}},v_{\mu,R_{j}},\tilde{U}_{\mu},\tilde{V}_{\mu})=0.$
@@ 8. The free boundary: regularity and structural properties
The remainder of the section focuses on the study of the singular part when $b\leq 0$. Given $X_{0}\in S(u)$, we use $N_{0}^{X_{0}}(r)$ to denote the Almgren frequency functional centered at $X_{0}$ (see (5.1)); to be precise,
(8.1) $N_{0}^{X_{0}}(r)\coloneqq N_{0}(r,u(X_{0}+\cdot),v(X_{0}+\cdot)).$
@@
With these definitions at hand, we proceed to prove that the blow-up limits $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ depend continuously on $X_{0}\in S_{\mu}(u)$.
@@
(8.4) $M_{\mu}^{\overline{X}}(r,u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})\leq M_{\mu}^{\overline{X}}(r_{\varepsilon},u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})+C(r_{\varepsilon}-r)\leq C\varepsilon$
holds for all $r\in(0,r_{\varepsilon})$. In particular, taking the limit as $r\to 0^{+}$ in the identity
$M_{\mu}^{\overline{X}}(r,u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})=\int_{S_{1}^{+}}y^{b}\left((u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{X_{0}})^{2}+(u_{\mu,r}^{\overline{X}}-\tilde{v}_{\mu}^{X_{0}})^{2}\right)\,d\mathcal{H}^{n},$
in view of (8.4) we obtain that
(8.5) $\int_{S_{1}^{+}}y^{b}\left((\tilde{u}_{\mu}^{\overline{X}}-\tilde{u}_{\mu}^{X_{0}})^{2}+(\tilde{v}_{\mu}^{\overline{X}}-\tilde{v}_{\mu}^{X_{0}})^{2}\right)\,d\mathcal{H}^{n}=\lim_{r\to 0^{+}}M_{\mu}^{\overline{X}}(r,u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})\leq C\varepsilon.$
@@
(8.6) $\int_{B_{1}^{+}}y^{b}\left((u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}})^{2}+(v_{\mu,r}^{\overline{X}}-\tilde{v}_{\mu}^{\overline{X}})^{2}\right)\,dX\leq C\varepsilon.$
@@
(8.7) $\leq\int_{0}^{1}2\rho^{n+b+2\mu}\left(M_{\mu}^{\overline{X}}(\rho r,u,v,\tilde{u}_{\mu}^{X_{0}},\tilde{v}_{\mu}^{X_{0}})+\int_{S_{1}^{+}}y^{b}\left((\tilde{u}_{\mu}^{\overline{X}}-\tilde{u}_{\mu}^{X_{0}})^{2}+(\tilde{v}_{\mu}^{\overline{X}}-\tilde{v}_{\mu}^{X_{0}})^{2}\right)\,d\mathcal{H}^{n}\right)\,d\rho.$
@@
(8.9) $\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{C^{0,\alpha}(B_{3/4}^{+})}\leq C\left(\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{L^{2}(B_{1}^{+};y^{b})}+\|r^{2}v_{\mu,r}^{\overline{X}}\|_{L^{q}(B_{1}^{+};y^{b})}\right).$
@@
Finally, if $n\geq 4$, since by assumption we have that $\mu<4/(n-3)$, then we can find $q\in((n+1)/2,2+2/\mu)$. This shows that there exists $C=C(\varepsilon)$, independent of $\overline{X}$ and $r$ such that $C(\varepsilon)\to 0^{+}$ as $\varepsilon\to 0^{+}$ and with the property that
(8.11) $\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{L^{\infty}(B_{3/4}^{+})}\leq C(\varepsilon)$
@@
(8.12) $\|v_{\mu,r}^{\overline{X}}-\tilde{v}_{\mu}^{\overline{X}}\|_{L^{\infty}(B_{1/2}^{+})}\leq C\left(\|v_{\mu,r}^{\overline{X}}-\tilde{v}_{\mu}^{\overline{X}}\|_{L^{2}(B_{3/4}^{+};y^{b})}+\|\bar{f}(\cdot,u_{\mu,r}^{\overline{X}})\|_{L^{\infty}(B_{3/4}^{\prime})}\right).$
@@
(8.14) $\|\bar{f}(\cdot,u_{\mu,r}^{\overline{X}})\|_{L^{\infty}(B_{3/4}^{\prime})}\leq Cr^{1-b+\mu(p-2)}\Big(\|u_{\mu,r}^{\overline{X}}-\tilde{u}_{\mu}^{\overline{X}}\|_{L^{\infty}(B_{3/4}^{\prime})}\\ +\|\tilde{u}_{\mu}^{\overline{X}}-\tilde{u}_{\mu}^{X_{0}}\|_{L^{\infty}(B_{3/4}^{\prime})}+\|\tilde{u}_{\mu}^{X_{0}}\|_{L^{\infty}(B_{3/4}^{\prime})}\Big)^{p-1}.$
@@
(8.15) $d_{\mu}^{X_{0}}\coloneqq\min\left\{d_{u,\mu}^{X_{0}},d_{v,\mu}^{X_{0}}\right\},$